[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Thu Jul 14, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu Jul 14 20:11:02 UTC 2016


HSU Shift Log for Thu Jul 14 2016 16:10:57 GMT-0400 (EDT)

Shift: H

Forecaster: R Berg

Technical Issues Summary:

ATCF issues: Craig and Monica (and others) have been working throughout the  
day to fix and determine the cause of the missing a-deck data for Celia  
(thanks and good luck!).

N-AWIPS/AWIPS2 issues:

Other Techincal issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160714/cf1234a6/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list